# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ III.1 Graphs and their associated matrices
Combinatorial graphs build up the mathematical foundations for complex networks. Here, we only define a graph, the graph Laplacian matrix, and the graph incidence matrix. The reader can refer to Section A.2 of the Appendix for other definitions in graph theory that are essential for this article.
@@
In this article, all the graphs have finite number of vertices. The vertex set of a graph $G$ can be written as $V(G)=\{v_{i}:i=1,2,\cdots,N\}$. Throughout the article, $N$ and $M$ denote the number of vertices and edges in the graph $G$, respectively. In the case of undirected graphs, with all of the edges being undirected, we consider only simple graphs (Definition A.2. 3). We define the Laplacian matrix of an undirected graph as follows:
@@
The Laplacian matrix $L(G)$ of a simple graph, or a weighted graph $G$, with $N$ vertices $v_{i}:i=1,2,\cdots,N$ is an $N\times N$ matrix L(G)=D(G)Q(G)L(G)=D(G)-Q(G) where $D(G)$ is the degree matrix (Definition A.2. 6) and $Q(G)$ is the adjacency matrix (Definition A.2. 7).
@@ III.2 Systems of linear equations arising from graphs
    As $G$ has $N$ vertices, the system size $\mathcal{N}=N$. $L$ is a square Hermitian matrix and it is positive semi-definite. At least one eigenvalue of $L$ is 0. As $L^{-1}$ does not exist, we compute $\vec{x}=L^{+}\vec{b}$, where $L^{+}$ is the pseudo-inverse of the matrix $L$ [penrose1956best].
@@
The system of linear equations arising from a graph Laplacian matrix is useful in electrical networks to find the effective resistance, $r_{\text{eff}}$, between two vertices [spielman2010algorithms]. Example 1 in Appendix A.3 illustrates it. Also, Example 2 in Appendix A.3 discusses a system of linear equations related to the incidence matrix of a graph that is used to model the traffic flow congestion.
@@ III.3 Categories of graph families
We define a graph family as a sequence of graphs $\mathcal{G}=\{G_{n}\}$, where $n$ is a natural number arranged in an increasing order.
The complexity of solving a NLSP using a QLS depends on the growth of $\mathcal{N}$, $\kappa(\mathcal{N})$ and $s(\mathcal{N})$. To understand the influence of these quantities on the runtime of a QLS, we numerically analyze them on different graph families. Based on the degree of quantum advantage, we classify graph families into different categories.
@@
A bad graph family is one for which the HHL algorithm yields no advantage relative to CLS.
Definition III.6.
A good graph family is one for which the HHL algorithm results in a sub-linear advantage relative to CLS.
Definition III.7.
A better graph family is one for which the HHL algorithm leads to a polynomial advantage relative to CLS.
Definition III.8.
A best graph family is one for which the HHL algorithm produces an exponential advantage relative to CLS.
Table 2: Table presenting our data on the Laplacian matrices of the considered graph families and their corresponding quantum advantage. The first column recalls our classification of graph families and their advantage with HHL over an efficient CLS: best (exponential), better (polynomial), good (sub-linear), and bad (no advantage). $\kappa$ and $s$ refer to the condition number and sparsity of a graph Laplacian of size $(N\times N)$, while $n\in\mathbb{Z}^{+}$. $t_{\mathrm{HHL}}(n)$ and $t_{\mathrm{CLS}}(n)$ denote the fitted runtime complexity expressions (excluding pre-factors) for HHL and CLS algorithms respectively. $\tilde{R}(n)=t_{\mathrm{CLS}}(n)/t_{\mathrm{HHL}}(n)$ decides the amount of advantage. $ll(n)$ is shorthand for $\mathrm{log}(\mathrm{log}(n))$. The last column indicates the advantage that we obtain for the same graph family with the CKS(1)/AQC(1) algorithm and the dream QLS, where ‘exp’, ‘poly’ and ‘sub-linear’ denote exponential, polynomial and sub-linear advantages respectively. ‘none’ means no advantage with the QLS.
Graph Graph name NN $\kappa$ $s$ $t_{\mathrm{HHL}}(n)$ $t_{\mathrm{CLS}}(n)$ Advantage with HHL, Advantage
type (Random (Yes/No)) with CKS(1)/AQC(1),
$\tilde{R}(n)=t_{\mathrm{CLS}}(n)/t_{\mathrm{HHL}}(n)$ with dream
solver
Best Hypercube graph (No) $2^{n}$ $n$ $n$ $n^{7}$ $2^{n}n^{3/2}\mathrm{log}(n)$ exp, $2^{n}\mathrm{log}(n)/n^{11/2}$ exp, exp
Better Modified Margulis- $n^{2}$ $\mathrm{log}^{2}(n)$ $c$ $\mathrm{log}^{8}(n)$ $n^{2}\mathrm{log}(n)ll(n)$ poly, $n^{2}ll(n)/\mathrm{log}^{7}(n)$ poly, poly
@@
Good Sudoku (No) $n^{4}$ $\mathrm{log}^{3}(n)$ $n^{2}$ $n^{4}\mathrm{log}^{11}(n)$ $n^{6}\mathrm{log}^{3/2}(n)ll(n)$ sub-linear, $n^{2}ll(n)/\mathrm{log}^{19/2}(n)$ poly, poly
Random regular $n$ $\mathrm{log}^{3}(n)$ $c$ $\mathrm{log}^{11}(n)$ $n\mathrm{log}^{3/2}(n)ll(n)$ sub-linear, $nll(n)/\mathrm{log}^{19/2}(n)$ sub-linear,
expander (Yes) sub-linear
Barabási-Albert (Yes) $n$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n\mathrm{log}^{9/2}(n)ll(n)$ sub-linear, $nll(n)/\mathrm{log}^{25/2}(n)$ sub-linear,
sub-linear
Newman-Watts $n$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{3}(n)$ $\mathrm{log}^{17}(n)$ $n\mathrm{log}^{9/2}(n)ll(n)$ sub-linear, $nll(n)/\mathrm{log}^{25/2}(n)$ sub-linear,
-Strogatz (Yes) sub-linear
Random regular (Yes) $n$ $\mathrm{log}^{2}(n)$ $c$ $\mathrm{log}^{8}(n)$ $n\mathrm{log}(n)ll(n)$ sub-linear, $nll(n)/\mathrm{log}^{7}(n)$ sub-linear,
sub-linear
Bad Grid 2d (No) $n$ $n^{2}$ $c$ $n^{6}\mathrm{log}^{2}(n)$ $n^{2}ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{2}(n)$ none, none
Hexagonal lattice (No) $n$ $n^{2}$ $c$ $n^{6}\mathrm{log}^{2}(n)$ $n^{2}ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{2}(n)$ none, none
Triangular lattice (No) $n$ $n$ $c$ $n^{3}\mathrm{log}^{2}(n)$ $n^{3/2}ll(n)$ none, $ll(n)/n^{3/2}\mathrm{log}^{2}(n)$ sub-linear,
sub-linear
Complete (No) $n$ $c$ $n$ $n^{2}\mathrm{log^{2}}(n)$ $n^{2}ll(n)$ none, $ll(n)/\mathrm{log}^{2}(n)$ sub-linear, poly
Turan (No) $n$ $\mathrm{log}^{2}(n)$ $n$ $n^{2}\mathrm{log}^{8}(n)$ $n^{2}\mathrm{log}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{7}(n)$ sub-linear, poly
Gaussian random $n$ $\mathrm{log}^{3}(n)$ $n$ $n^{2}\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{19/2}(n)$ sub-linear,
partition (Yes) poly
Geographical threshold (Yes) $n$ $\mathrm{log}^{3}(n)$ $n$ $n^{2}\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{19/2}(n)$ sub-linear, poly
Soft random geometric (Yes) $n$ $\mathrm{log}^{3}(n)$ $n$ $n^{2}\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{19/2}(n)$ sub-linear, poly
Thresholded random $n$ $\mathrm{log}^{3}(n)$ $n$ $n^{2}\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{19/2}(n)$ sub-linear,
geometric (Yes) poly
Planted partition (Yes) $n$ $\mathrm{log}^{3}(n)$ $n$ $n^{2}\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{19/2}(n)$ sub-linear, poly
Random geometric (Yes) $n$ $\mathrm{log}^{3}(n)$ $n$ $n^{2}\mathrm{log}^{11}(n)$ $n^{2}\mathrm{log}^{3/2}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{19/2}(n)$ sub-linear, poly
Uniform random $n$ $c$ $n$ $n^{2}\mathrm{log^{2}}(n)$ $n^{2}ll(n)$ none, $ll(n)/\mathrm{log}^{2}(n)$ sub-linear,
intersection (Yes) poly
Hk,n Harary (No) $n$ $n^{2}$ $c$ $n^{6}\mathrm{log}^{2}(n)$ $n^{2}ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{2}(n)$ none, none
Hm,n Harary (No) $n$ $n^{2}$ $c$ $n^{6}\mathrm{log}^{2}(n)$ $n^{2}ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{2}(n)$ none, none
Circular ladder (No) $n$ $n^{2}$ $c$ $n^{6}\mathrm{log}^{2}(n)$ $n^{2}ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{2}(n)$ none, none
Ladder (No) $n$ $n^{2}$ $c$ $n^{6}\mathrm{log}^{2}(n)$ $n^{2}ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{2}(n)$ none, none
Ring of cliques (No) $n$ $n^{2}$ $c$ $n^{6}\mathrm{log}^{2}(n)$ $n^{2}ll(n)$ none, $ll(n)/n^{4}\mathrm{log}^{2}(n)$ none, none
Balanced binary tree (No) $2^{n}$ $2^{n}$ $c$ $n^{2}2^{3n}$ $2^{3n/2}\mathrm{log}(n)$ none, $\mathrm{log}(n)/n^{2}2^{3n/2}$ exp, exp
Balanced ternary tree (No) $3^{n}$ $3^{n}$ $c$ $n^{2}3^{3n}$ $3^{3n/2}\mathrm{log}(n)$ none, $\mathrm{log}(n)/n^{2}3^{3n/2}$ exp, exp
Binomial tree (No) $2^{n}$ $2^{n}$ $n$ $n^{4}2^{3n}$ $2^{3n/2}n\mathrm{log}(n)$ none, $\mathrm{log}(n)/n^{3}2^{3n/2}$ exp, exp
Grid 2d graph ($r=c$) (No) $4^{n}$ $4^{n}$ $c$ $n^{2}4^{3n}$ $4^{3n/2}\mathrm{log}(n)$ none, $\mathrm{log}(n)/n^{2}4^{3n/2}$ exp, exp
Random lobster (Yes) $n$ $n^{2}$ $n^{2}$ $n^{10}\mathrm{log}^{2}(n)$ $n^{4}ll(n)$ none, $ll(n)/n^{6}\mathrm{log}^{2}(n)$ none, sub-linear
Gn,p random (Yes) $n$ $\mathrm{log}^{2}(n)$ $n$ $n^{2}\mathrm{log}^{8}(n)$ $n^{2}\mathrm{log}(n)ll(n)$ none, $ll(n)/\mathrm{log}^{7}(n)$ sub-linear,
poly
Before we proceed to the results, we summarize our typical workflow for each graph that we pick. As outlined in the earlier sections, we start with a graph and a particular construction for generating a graph family. We plot $\kappa(\mathcal{N})$ of Laplacian or incidence matrix as well as $s(\mathcal{N})$ against system size, $\mathcal{N}$. We then fit the data points to appropriate functions, and plot $R(\mathcal{N})$ versus $\mathcal{N}$ to check if a curve goes above $1$ within our dataset. We recall that $R(\mathcal{N})>1$ is the regime of quantum advantage. We also check the system size growth, $\mathcal{N}$ versus $n$, and assuming that the function fits for $\kappa(\mathcal{N})$ and $s(\mathcal{N})$ hold for much larger system sizes. We check how much advantage we can expect from the graph family, that is, assess to which of the aforementioned set of categories our graph family belongs: bad, good, better, or best. All the graph generation steps were carried out in NetworkX (version 3.4.2) [networkx]. We also add that for brevity, we hereafter call $\kappa(\mathcal{N})$, for example, as just $\kappa$, and so on.
@@ IV Results from our survey of networks
In this section, we present the results for Laplacian (Sections IV.1) and incidence matrix-based graph families (Section IV.2). We note that our numerical study on graph families involves a number of assumptions, which in turn set the scope of our study. Section A.4 of the Appendix extensively discusses the assumptions and scope of this study, and the discussion is supported by error analysis on certain graph families. Section A.5 of the Appendix provides the details of all of the input parameters along with the construction for the graph families considered. In the subsequent sub-sections, we provide for each graph family a figure that contains three sub-figures:
@@
  • The second sub-figure presents the runtime ratio, $R(\mathcal{N})$, between each of the considered QLSs and the CLS. It also highlights the crossover at $R(\mathcal{N})=1$, above which the QLS outperforms CLS, even within the limited range of $\mathcal{N}$ values considered in our datasets.
  • The third sub-figure shows the degree of quantum advantage achieved by HHL in terms of $\tilde{R}(n)$, which in turn is used to categorize the graph family as best/better/good/bad. A reference line $\tilde{R}(n)=n$ is included to visually identify the cases exhibiting a sub-linear advantage. In this sub-figure, all pre-factors from our fits are excluded, as we are interested in very large $\mathcal{N}$ behaviour by assuming that our fit functions remain valid at those regimes.
(a) (b) (c)
Figure 2: Data for the hypercube graph family (best graph family) from Laplacian matrix based systems: sub-figures showing (a) $\kappa$ and $s$ versus $N$, (b) $R$ versus $N$, as well as (c) $\tilde{R}$ versus $n$, and a reference linear curve, $n$.
(a) (b) (c)
Figure 3: Better graph family from Laplacian matrix based system: sub-figures showing (a) $\kappa$ and $s$ versus $N$, (b) $R$ versus $N$, as well as (c) $\tilde{R}$ versus $n$ and a reference linear curve, $n$, for the modified Margulis-Gabber-Galil graph.
(a) (b) (c)
Figure 4: Best graph family from incidence-matrix based system: sub-figures showing (a) $\kappa$ and $s$ versus $N^{\prime}$, (b) $R$ versus $N^{\prime}$, as well as (c) $\tilde{R}$ versus $n$ and a reference linear curve, $n$, for the directed hypercube graph.

### v4

@@ III.1 Graphs and their associated matrices
Combinatorial graphs build up the mathematical foundations for complex networks. Here, we only define a graph, the graph Laplacian matrix, and the graph incidence matrix. The reader can refer to Section S.2 of the Supplemental Material for other definitions in graph theory that are essential for this work.
@@
In this article, all the graphs have finite number of vertices. The vertex set of a graph $G$ can be written as $V(G)=\{v_{i}:i=1,2,\cdots,N\}$. Throughout the article, $N$ and $M$ denote the number of vertices and edges in the graph $G$, respectively. In the case of undirected graphs, with all of the edges being undirected, we consider only simple graphs. We define the Laplacian matrix of an undirected graph as follows:
@@
The Laplacian matrix $L(G)$ of a simple graph, or a weighted graph $G$, with $N$ vertices $v_{i}:i=1,2,\cdots,N$ is an $N\times N$ matrix L(G)=D(G)Q(G)L(G)=D(G)-Q(G), where $D(G)$ is the degree matrix and $Q(G)$ is the adjacency matrix.
@@ III.2 Systems of linear equations arising from graphs
    As $G$ has $N$ vertices, the system size $\mathcal{N}=N$. $L$ is a square Hermitian matrix and it is positive semi-definite. At least one eigenvalue of $L$ is 0. As $L^{-1}$ does not exist, we compute $\vec{x}=L^{+}\vec{b}$, where $L^{+}$ is the pseudo-inverse of the matrix $L$ [53].
@@
The system of linear equations arising from a graph Laplacian matrix is useful in electrical networks to find the effective resistance, $r_{\text{eff}}$, between two vertices [60]. Example A of Section S.3 in the Supplemental Material illustrates it. Also, Example B of the same section discusses a system of linear equations related to the incidence matrix of a graph that is used to model the traffic flow congestion.
@@ III.3 Categories of graph families
The complexity of solving a NLSP using a QLS depends on the growth of $\kappa(\mathcal{N})$ and $s(\mathcal{N})$. To understand the influence of these quantities on the runtime of a QLS, we numerically analyze them on different graph families. As HHL algorithm is the prototypical QLS, with the other QLSs considered in this study improving upon its runtime complexity, we classify the graph families into two categories based on whether HHL offers advantage or not. The categories are defined below:
@@
A good graph family is one for which the HHL algorithm gives advantage in its runtime relative to CLS. Here, advantage could be an exponential or a polynomial separation between the runtimes of HHL algorithm and CLS.
@@
A bad graph family is one for which the HHL algorithm yields no advantage with respect to CLS.
If the HHL algorithm offers an advantage for some graph family, other QLSs considered in the study will too. However, there could be graph families where HHL does not offer an advantage, but improvements over it could.
Before we proceed to the results, we summarize our typical workflow for each graph family that we pick. We start with a graph and a particular construction for generating a graph family. We plot $\kappa(\mathcal{N})$ of Laplacian or incidence matrix as well as $s(\mathcal{N})$ against system size, $\mathcal{N}$. We then fit the data points to appropriate functions, and plot $R(\mathcal{N})$ versus $\mathcal{N}$ to check if a curve goes above $1$ within our dataset. We recall that $R(\mathcal{N})>1$ is the regime of quantum advantage. We assume that our function fits for $\kappa(\mathcal{N})$ and $s(\mathcal{N})$ hold for much larger system sizes. We check if advantage can be expected from the graph family, that is, assess to which of the aforementioned set of categories our graph family belongs: good or bad. All the graph generation steps were carried out in NetworkX (version 3.4.2) [28]. For brevity, we hereafter call $\kappa(\mathcal{N})$, for example, as just $\kappa$, and so on.
We add that while the interplay between $\kappa$ and $s$ scaling as functions of $\mathcal{N}$ decide the amount of advantage, the allowable values of $\mathcal{N}$ for a given graph family decides its classical hardness. Depending on the graph construction, $\mathcal{N}=f(n)$, where $n\in\mathbb{N}$.
[…]
Graph name $\kappa$ $s$ $t_{\mathrm{HHL}}$ $t_{\mathrm{CKS(1)}}(\text{or }t_{\mathrm{AQC(1)}})$ $t_{\mathrm{dream\ QLS}}$ $t_{\mathrm{CLS}}$
(Random (Yes/No))
Hypercube graph (No) $\mathrm{log}(N)$ $\mathrm{log}(N)$ $\mathrm{log}^{7}(N)$ $\mathrm{log}^{3}(N)ll(N)$ $\mathrm{log}^{5/2}(N)ll(N)$ $N\mathrm{log}^{3/2}(N)ll(N)$
Modified Margulis- $\mathrm{log}^{2}(N)$ $c$ $\mathrm{log}^{8}(N)$ $\mathrm{log}^{3}Nll(N)$ $\mathrm{log}^{3}(N)ll(N)$ $N\mathrm{log}(N)ll(N)$
@@
Sudoku (No) $\mathrm{log}^{3}(N)$ $\sqrt{N}$ $N\mathrm{log}^{11}(N)$ $\sqrt{N}\mathrm{log}^{4}(N)\mathrm{log}(\sqrt{N}\mathrm{log}^{4}(N))$ $N^{1/4}\mathrm{log}^{4}(N)ll(N)$ $N^{3/2}\mathrm{log}^{3/2}(N)ll(N)$
Barabási-Albert (Yes) $\mathrm{log}^{3}(N)$ $\mathrm{log}^{3}(N)$ $\mathrm{log}^{17}(N)$ $\mathrm{log}^{7}(N)ll(N)$, $\mathrm{log}^{11/2}(N)ll(N)$ $N\mathrm{log}^{9/2}(N)ll(N)$
Newman-Watts $\mathrm{log}^{3}(N)$ $\mathrm{log}^{3}(N)$ $\mathrm{log}^{17}(N)$ $\mathrm{log}^{7}(N)ll(N)$ $\mathrm{log}^{11/2}(N)ll(N)$ $N\mathrm{log}^{9/2}(N)ll(N)$
-Strogatz (Yes)
Random regular $\mathrm{log}^{3}(N)$ $c$ $\mathrm{log}^{11}(N)$ $\mathrm{log}^{4}(N)ll(N)$ $\mathrm{log}^{4}(N)ll(N)$ $N\mathrm{log}^{3/2}(N)ll(N)$
expander (Yes)
Random regular (Yes) $\mathrm{log}^{2}(N)$ $c$ $\mathrm{log}^{8}(N)$ $\mathrm{log}^{3}(N)ll(N)$ $\mathrm{log}^{3}(N)ll(N)$ $N\mathrm{log}(N)ll(N)$
@@ IV Results from our survey of networks
Table 3: Table presenting our data for Laplacian matrix-based bad graph families that offer no advantage with the HHL algorithm. We also give the advantage offered by these graph families with CKS(1) (or AQC(1)) (see Eq. 2) and dream QLS algorithms (see Eq. 3). ‘Nil’ represents no advantage and ‘Exp’ refers to exponential advantage in the table.
Graph name $\kappa$ $s$ $t_{\mathrm{HHL}}$ $t_{\mathrm{CKS(1)}}(\text{or }t_{\mathrm{AQC(1)}}),$ $t_{\mathrm{dream\ QLS}}$ $t_{\mathrm{CLS}}$ Advantage with
(Random (Yes/No)) HHL,
CKS(1)(or AQC(1)),
dream QLS
Grid 2d (No) $N^{2}$ $c$ $N^{6}\mathrm{log}^{2}(N)$ $N^{2}\mathrm{log}(N)\mathrm{log}(N^{2}\mathrm{log}(N))$ $N^{2}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Nil, Nil
Hexagonal $N^{2}$ $c$ $N^{6}\mathrm{log}^{2}(N)$ $N^{2}\mathrm{log}(N)\mathrm{log}(N^{2}\mathrm{log}(N))$, $N^{2}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Nil, Nil
lattice (No)
Triangular NN $c$ $N^{3}\mathrm{log}^{2}(N)$ $N\mathrm{log}(N)\mathrm{log}(N\mathrm{log}(N))$ $N\mathrm{log}(N)ll(N)$ $N^{3/2}ll(N)$ Nil, Exp, Exp
lattice (No)
Complete (No) $c$ NN $N^{2}\mathrm{log}^{2}(N)$ $N\mathrm{log}(N)\mathrm{log}(N\mathrm{log}(N))$ $\sqrt{N}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Exp, Exp
Turan (No) $\mathrm{log}^{2}(N)$ NN $N^{2}\mathrm{log}^{8}(N)$ $N\mathrm{log}^{3}(N)\mathrm{log}(N\mathrm{log}^{3}(N))$ $\sqrt{N}\mathrm{log}^{3}(N)ll(N)$ $N^{2}\mathrm{log}(N)ll(N)$ Nil, Exp, Exp
Gaussian random $\mathrm{log}^{3}(N)$ NN $N^{2}\mathrm{log}^{11}(N)$ $N\mathrm{log}^{4}(N)\mathrm{log}(N\mathrm{log}^{4}(N))$ $\sqrt{N}\mathrm{log}^{4}(N)ll(N)$ $N^{2}\mathrm{log}^{3/2}(N)ll(N)$ Nil, Exp, Exp
partition (Yes)
Geographical $\mathrm{log}^{3}(N)$ NN $N^{2}\mathrm{log}^{11}(N)$ $N\mathrm{log}^{4}(N)\mathrm{log}(N\mathrm{log}^{4}(N))$ $\sqrt{N}\mathrm{log}^{4}(N)ll(N)$ $N^{2}\mathrm{log}^{3/2}(N)ll(N)$ Nil, Exp, Exp
threshold (Yes)
Soft random $\mathrm{log}^{3}(N)$ NN $N^{2}\mathrm{log}^{11}(N)$ $N\mathrm{log}^{4}(N)\mathrm{log}(N\mathrm{log}^{4}(N))$ $\sqrt{N}\mathrm{log}^{4}(N)ll(N)$ $N^{2}\mathrm{log}^{3/2}(N)ll(N)$ Nil, Exp, Exp
geometric (Yes)
Thresholded random $\mathrm{log}^{3}(N)$ NN $N^{2}\mathrm{log}^{11}(N)$ $N\mathrm{log}^{4}(N)\mathrm{log}(N\mathrm{log}^{4}(N))$ $\sqrt{N}\mathrm{log}^{4}(N)ll(N)$ $N^{2}\mathrm{log}^{3/2}(N)ll(N)$ Nil, Exp, Exp
geometric (Yes)
Planted $\mathrm{log}^{3}(N)$ NN $N^{2}\mathrm{log}^{11}(N)$ $N\mathrm{log}^{4}(N)\mathrm{log}(N\mathrm{log}^{4}(N))$ $\sqrt{N}\mathrm{log}^{4}(N)ll(N)$ $N^{2}\mathrm{log}^{3/2}(N)ll(N)$ Nil, Exp, Exp
partition (Yes)
Random $\mathrm{log}^{3}(N)$ NN $N^{2}\mathrm{log}^{11}(N)$ $N\mathrm{log}^{4}(N)\mathrm{log}(N\mathrm{log}^{4}(N))$ $\sqrt{N}\mathrm{log}^{4}(N)ll(N)$ $N^{2}\mathrm{log}^{3/2}(N)ll(N)$ Nil, Exp, Exp
geometric (Yes)
Uniform random $c$ NN $N^{2}\mathrm{log}^{2}(N)$ $N\mathrm{log}(N)\mathrm{log}(N\mathrm{log}(N))$ $\sqrt{N}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Exp, Exp
intersection (Yes)
Hk,n Harary (No) $N^{2}$ $c$ $N^{6}\mathrm{log}^{2}(N)$ $N^{2}\mathrm{log}(N)\mathrm{log}(N^{2}\mathrm{log}(N))$ $N^{2}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Nil, Nil
Hn,m Harary (No) $N^{2}$ $c$ $N^{6}\mathrm{log}^{2}(N)$ $N^{2}\mathrm{log}(N)\mathrm{log}(N^{2}\mathrm{log}(N))$ $N^{2}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Nil, Nil
Circular ladder (No) $N^{2}$ $c$ $N^{6}\mathrm{log}^{2}(N)$ $N^{2}\mathrm{log}(N)\mathrm{log}(N^{2}\mathrm{log}(N))$ $N^{2}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Nil, Nil
Ladder (No) $N^{2}$ $c$ $N^{6}\mathrm{log}^{2}(N)$ $N^{2}\mathrm{log}(N)\mathrm{log}(N^{2}\mathrm{log}(N))$ $N^{2}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Nil, Nil
Ring of cliques (No) $N^{2}$ $c$ $N^{6}\mathrm{log}^{2}(N)$ $N^{2}\mathrm{log}(N)\mathrm{log}(N^{2}\mathrm{log}(N))$ $N^{2}\mathrm{log}(N)ll(N)$ $N^{2}ll(N)$ Nil, Nil, Nil
Balanced binary NN $c$ $N^{3}\mathrm{log}^{2}(N)$ $N\mathrm{log}(N)\mathrm{log}(N\mathrm{log}(N))$ $N\mathrm{log}(N)ll(N)$ $N^{3/2}ll(N)$ Nil, Exp, Exp
tree (No)
Balanced ternary NN $c$ $N^{3}\mathrm{log}^{2}(N)$ $N\mathrm{log}(N)\mathrm{log}(N\mathrm{log}(N))$ $N\mathrm{log}(N)ll(N)$ $N^{3/2}ll(N)$ Nil, Exp, Exp
tree (No)
Binomial tree (No) NN $\mathrm{log}(N)$ $N^{3}\mathrm{log}^{4}(N)$ $N\mathrm{log}^{2}(N)\mathrm{log}(N\mathrm{log}^{2}(N))$ $N\mathrm{log}^{3/2}(N)ll(N)$ $N^{3/2}\mathrm{log}(N)ll(N)$ Nil, Exp, Exp
Grid 2d NN $c$ $N^{3}\mathrm{log}^{2}(N)$ $N\mathrm{log}(N)\mathrm{log}(N\mathrm{log}(N))$ $N\mathrm{log}(N)ll(N)$ $N^{3/2}ll(N)$ Nil, Exp, Exp
graph ($r=c$) (No)
Random $N^{2}$ $N^{2}$ $N^{10}\mathrm{log}^{2}(N)$ $N^{4}\mathrm{log}(N)\mathrm{log}(N^{4}\mathrm{log}(N))$ $N^{3}\mathrm{log}(N)ll(N)$ $N^{4}ll(N)$ Nil, Nil, Exp
lobster (Yes)
Gn,p random (Yes) $\mathrm{log}^{2}(N)$ NN $N^{2}\mathrm{log}^{8}(N)$ $N\mathrm{log}^{3}(N)\mathrm{log}(N\mathrm{log}^{3}(N))$ $\sqrt{N}\mathrm{log}^{3}(N)ll(N)$ $N^{2}\mathrm{log}(N)ll(N)$ Nil, Exp, Exp
@@
In this section, we present the results for Laplacian (Sections IV.1) and incidence matrix-based graph families (Section IV.2). We note that our numerical study on graph families involves a number of assumptions, which in turn set the scope of our study. Section S.4 of the Supplemental Material extensively discusses the assumptions and scope of this study, and the discussion is supported by error analysis on certain graph families. Section S.5 of the Supplemental Material provides the details on all of the input parameters along with the construction for the graph families considered. We try to ensure diversity in terms of $\kappa$, $s$ and system size growth when we pick the graph families. Hence, we have considered non-random graphs, random graphs, trees, and expanders in our graph families. We note at this juncture that to the best of our knowledge, among the 50 graph families that we pick, only the hypercube graph and the complete graph have known behaviours for their $\kappa$ scaling. In the subsequent sub-sections, we provide for each graph family a figure that contains two sub-figures:
@@
  • The second sub-figure presents the runtime ratio, $R(\mathcal{N})$, between each of the considered QLSs (listed in Table 1 and the CLS. It also highlights the crossover at $R(\mathcal{N})=1$, above which the QLS outperforms CLS, even within the limited range of $\mathcal{N}$ values considered in our datasets.
Furthermore, we present the runtime expressions for HHL, CKS(1) (whose runtime complexity is the same as AQC(1)), dream QLS and CLS algorithms for all graph families considered in our study in Tables 2, 3, 4 and 5. In these tables, we also compare runtime advantage from HHL which is the prototypical QLS, CKS(1) (or AQC(1)) which has near optimal runtime complexity among the QLSs considered in the study, and the dream QLS which is optimal in its runtime complexity.
(a) (b)
(c) (d)
Figure 2: The figure presents two instances of good (advantage with HHL) graph families for Laplacian matrix-based systems. The first sub-figure of each panel shows condition number, $\kappa$, and sparsity, $s$, versus the number of vertices, $N$, and second sub-figure displays the runtime ratio, $R(N)$, versus $N$ for every QLS considered in the study.
(a) (b)
(c) (d)
Figure 3: The figure illustrates two instances of bad (no advantage with HHL) graph families for Laplacian matrix-based systems. The first sub-figure of each panel represents growth of condition number, $\kappa$, and sparsity, $s$, with number of vertices NN while the second sub-figure shows the runtime ratio $R(N)$ with NN for every QLS considered in the study.
(a) (b)
(c) (d)
(e) (f)
(g) (h)
Figure 4: The figure presents two instances each for good and bad categories of graph families for incidence matrix-based systems. Sub-figures (a)-(d) belong to good graph category, whereas (e)-(h) belong to bad graph category.